An order is written for 10 mL of a 10% calcium chloride injection and 10 mL of multivitamin injection (MVI) to be added to 50o mL of D5W. The infusion is to be administered over 6 hours. The Ⅳset delivers 15 drops/ mL. what should be the rate of flow in drops per minute to deliver this infusion? 7.
Answer #1

Ans) Dose: 10 ml of 10% calcium chloride injection + 10 ml of inj MVI in 500ml of D5W.= 520 ml

Time duration: over 6hrs= 6× 60= 360 mins

Drop factor: 15 drops/min

Rate of flow in drop/min= Total amount to be infused in ml÷ time duration in mins × drop factor

= 520÷ 360 ×15

= 21.66 drops/min

The rate of flow is 21.66 drops/min.